If you hooked two springs together end to end and then suspended a 200g mass from the end, what would the overall elongation be? Therefore, what would be the effective spring constant of the two springs considered as one?

Related equations from the lab:

F = - kx

T = 2π * (m/k)^(1/2)

Mg = kx
Answer #1

since the spring are connedted to end to end it is series connection

1/ k eff = 1/k+ 1/k

keff = k/2

Apply Hooks law

F = k eff x

the overall elongation is

x= F/keff = mg/k/2 = 2 mg/k = 2 (0.2)(9.8)/k = 3.92 /k

the overall elongation is x= 3.92/k